# Deep FlexQP: Accelerated Nonlinear Programming via Deep Unfolding

### v1

@@ 3.3 Operator Splitting & ADMM
The optimization in Equation 6 can be simplified further to make the terms appearing in the $\ell_{1}$ penalty functions easier to handle. Introducing decision variables $z_{I}\in{\mathbb{R}}^{m}$ and $z_{E}\in{\mathbb{R}}^{p}$, we have
$\begin{aligned} \operatorname*{minimize}_{x,s,z_{I},z_{E}}\quad&\frac{1}{2}x^{\top}Px+q^{\top}x+\mu_{I}\left\lVert z_{I}\right\rVert_{1}+\mu_{E}\left\lVert z_{E}\right\rVert_{1},\\ \operatorname*{subject\ to}\quad&z_{I}=Gx+s-h,\\ &z_{E}=Ax-b,\\ &s\geq 0.\end{aligned}$ (7)
This is analogous to the transformation used by OSQP (stellato2020osqp) and will help simplify the ADMM updates. As the variables $z_{I}$ and $z_{E}$ are equal to the constraint violation, their optimal values can be viewed as a certificate of feasibility for Equation 3 if $z_{I}^{*}=z_{E}^{*}=0$ and infeasibility if $z_{I}^{*}\neq 0$ or $z_{E}^{*}\neq 0$. While it may seem tempting to apply ADMM to this formulation, the resultant updates will not have a closed-form solution no matter how the variable splitting is performed. Therefore, we perform a final transformation by introducing copy variables $\tilde{x}$ $\tilde{s}$, $\tilde{z}_{I}$, and $\tilde{z}_{E}$, yielding
$\begin{aligned} \operatorname*{minimize}_{\tilde{\bm{x}},\bm{x}}\quad&\underbrace{\frac{1}{2}\tilde{x}^{\top}P\tilde{x}+q^{\top}\tilde{x}+{\mathcal{I}}_{I}(\tilde{\bm{x}})+{\mathcal{I}}_{E}(\tilde{\bm{x}})}_{=\vcentcolon\,f(\tilde{\bm{x}})}+\underbrace{{\mathcal{I}}_{s}(\bm{x})+\mu_{I}\left\lVert z_{I}\right\rVert_{1}+\mu_{E}\left\lVert z_{E}\right\rVert_{1}}_{=\vcentcolon\,g(\bm{x})},\\ \operatorname*{subject\ to}\quad&\tilde{\bm{x}}=\bm{x},\end{aligned}$ (8)
@@
${\mathcal{I}}_{I}(\bm{x})=\begin{cases}0&z_{I}=Gx+s-h,\\ +\infty&\text{otherwise},\end{cases}\quad{\mathcal{I}}_{E}(\bm{x})=\begin{cases}0&z_{E}=Ax-b,\\ +\infty&\text{otherwise},\end{cases}$
${\mathcal{I}}_{s}(\bm{x})=\begin{cases}0&s\geq 0,\\ +\infty&\text{otherwise}.\end{cases}$
Let the dual variables for the constraint $\tilde{\bm{x}}=\bm{x}$ be $\bm{y}=(w_{x},w_{s},y_{I},y_{E})$. The ADMM updates for solving Equation 8 are given by
$\begin{aligned} \tilde{\bm{x}}^{k+1}=\operatorname*{arg\,min}_{\tilde{\bm{x}}}\hskip 5.0ptf(\tilde{\bm{x}})+(\sigma_{x}/2)\lVert\tilde{x}-x^{k}+\sigma_{x}^{-1}w_{x}^{k}\rVert_{2}^{2}+(\sigma_{s}/2)\lVert\tilde{s}-s^{k}+\sigma_{s}^{-1}w_{s}^{k}\rVert_{2}^{2},\\ +(\rho_{I}/2)\lVert\tilde{z}_{I}-z_{I}^{k}+\rho_{I}^{-1}y_{I}^{k}\rVert_{2}^{2}+(\rho_{E}/2)\lVert\tilde{z}_{E}-z_{E}^{k}+\rho_{E}^{-1}y_{E}^{k}\rVert_{2}^{2}\end{aligned}$ (9a)
$x^{k+1}$ $=\alpha\tilde{x}^{k+1}+(1-\alpha)x^{k}+\sigma_{x}^{-1}w_{x}^{k},$ (9b)
$s^{k+1}$ $=\big(\alpha\tilde{s}^{k+1}+(1-\alpha)s^{k}+\sigma_{s}^{-1}w_{s}^{k}\big)_{+},$ (9c)
$z_{I}^{k+1}$ $=S_{\mu_{I}/\rho_{I}}\big(\alpha\tilde{z}_{I}^{k+1}+(1-\alpha)z_{I}^{k}+\rho_{I}^{-1}y_{I}^{k}\big),$ (9d)
$z_{E}^{k+1}$ $=S_{\mu_{E}/\rho_{E}}\big(\alpha\tilde{z}_{E}^{k+1}+(1-\alpha)z_{E}^{k}+\rho_{E}^{-1}y_{E}^{k}\big),$ (9e)
$w_{x}^{k+1}$ $=w_{x}^{k}+\sigma_{x}(\tilde{x}^{k+1}-x^{k+1}),$ (9f)
$w_{s}^{k+1}$ $=w_{s}^{k}+\sigma_{s}(\tilde{s}^{k+1}-s^{k+1}),$ (9g)
$y_{I}^{k+1}$ $=y_{I}^{k}+\rho_{I}(\tilde{z}_{I}^{k+1}-z_{I}^{k+1}),$ (9h)
$y_{E}^{k+1}$ $=y_{E}^{k}+\rho_{E}(\tilde{z}_{E}^{k+1}-z_{E}^{k+1}),$ (9i)
where $\sigma_{x},\sigma_{s},\rho_{I},\rho_{E}>0$ are the augmented Lagrangian penalty parameters, $\alpha\in(0,2)$ is the ADMM relaxation parameter, $(s)_{+}=\max(s,0)$ is the rectified linear unit (ReLU) activation function, and $S_{\kappa}(z)=(z-\kappa)_{+}-(-z-\kappa)_{+}$ is the soft thresholding operator, which is the proximal operator of the $\ell_{1}$ norm (boyd2011distributed). Note that by Equations 9b and 9f, we have that $w_{x}^{k+1}=0$ for all $k\geq 0$, so the $w_{x}$ variable and update can be disregarded. The first block update Equation 9a is the most computationally-demanding step of the algorithm and requires the solution of an equality-constrained QP. We show how to solve this QP using either a direct or indirect method in Appendix B; the indirect method becomes the only suitable choice for large-scale problems where the dimension can be very large. The final algorithm is summarized in Algorithm 1 of Appendix C.
@@ 3.4 Convergence & Theoretical Analysis
We establish the convergence of Algorithm 1 by showing that there always exists a saddle point of the Lagrangian for Equation 8 for any $\mu_{I},\mu_{E}>0$, as long as the relaxed QP objective in Equation 6 is not unbounded below. Then, since the optimization is a composite minimization of two closed, proper, convex functions, the algorithm converges by the general convergence of two-block ADMM (boyd2011distributed). The proof is provided in Appendix D.
@@
$\mathcal{L}(\tilde{\bm{x}},\bm{x},\bm{y})=f(\tilde{\bm{x}})+g(\bm{x})+w_{x}^{\top}(\tilde{x}-x)+w_{s}^{\top}(\tilde{s}-s)+y_{I}^{\top}(\tilde{z}_{I}-z_{I})+y_{E}^{\top}(\tilde{z}_{E}-z_{E}).$ (10)
The assumption on the objective being unbounded below is called coercitivity and is relatively weak (bauschke2017convex). Coercitivity is only broken in the rare case when there exists an $x\neq 0$ such that $Px=0$, $Gx=0$, $Ax=0$, and $q^{\top}x<0$, which causes the optimal solution to diverge. In our cases of interest, we consider over-constrained problems with many more constraints than optimization variables, thus it is extremely unlikely that this assumption does not hold. Moreover, a bounded objective can be guaranteed if $P\succ 0$ or if $G$ or $A$ are full column rank.
The following theorem establishes the relationship between the FlexQP solution and the solution to the original QP and can be proven using the definition of soft thresholding. The proof is given in Appendix E.
@@
This shows that FlexQP solves Equation 3 if the original QP was feasible, and otherwise identifies a stationary point of the infeasibilities, similar to nocedal2006numerical.
@@
Role of Elastic Penalty Parameters: The parameters $\mu_{I}$ and $\mu_{E}$ only appear in a single step of the algorithm during Equations 9d and 9e as part of the soft thresholding in the second block ADMM updates. Larger elastic penalties $\mu$ result in a larger threshold, meaning that a larger amount of constraint violation will be zeroed out. The choice of $\mu$ is key for satisfying the conditions of Theorem 3.1.
@@
The parameter $\sigma_{s}$ plays the role of quadratic cost on the slack variable $s$ when solving Equation 9a. It also plays a small role in regularizing the constraint matrix $G$. In practice, tuning this parameter is the most difficult as the optimal value appears to depend strongly on the scaling of the objective and the constraints. This motivates our adaptive data-driven approach described in Section 4.
The penalty parameters $\rho_{I}$ and $\rho_{E}$ play two key roles in the algorithm. First, they regularize the constraint matrices $G$ and $A$ so that Equation 9a is solvable regardless of the rank of $G$ or $A$. Second, they weight the noise level in the soft thresholding operations Equations 9d and 9e playing an inverse role to $\mu_{I}$ and $\mu_{E}$, where a larger $\rho$ results in a smaller threshold. Determining the optimal values of these parameters by hand is unintuitive as they can have varying effects on the optimization, further motivating the deep unfolding approach presented in the next section.
@@ 4 Accelerating Quadratic Programming via Deep Unfolding
Figure 2: One layer of our proposed Deep FlexQP architecture. We learn dimension-agnostic feedback policies for the parameters while the propagation from one layer to the next is defined by the ADMM updates Equation 9.
We focus our study on two recently proposed data-accelerated QP optimizers. The deep centralized QP optimizer from saravanos2025deep is a version of deep-unfolded OSQP where the penalty parameters $\rho$ and relaxation parameter $\alpha$ are learned as feedback policies on the problem residuals using an analogy to feedback control. In our comparisons, we refer to their method as Deep OSQP. The main limitation of their approach is that only scalar penalty parameters are learned, but it could be the case that different penalty parameters should be applied to different constraints to more effectively accelerate the optimizer. This was the main motivation for the deep QP method proposed by ichnowski2021accelerating, where a policy that outputs a vector of penalty parameters is learned using reinforcement learning. The vector policy is applied across the constraint dimensions so it is dimension-agnostic and generalizes across different problem classes. The authors show that the vector policy outperforms the scalar one across a suite of QP benchmarks. However, unlike saravanos2025deep, the authors do not learn the relaxation parameter $\alpha$, which can greatly improve the convergence of ADMM (boyd2011distributed). We implement the approach from ichnowski2021accelerating and train it using the supervised learning scheme from saravanos2025deep, leading to the baseline Deep OSQP — RLQP Parameterization. Finally, we implement a best-of-both-worlds approach that learns a vector feedback policy for the penalty parameters $\rho$ while also learning a policy for the ADMM relaxation parameter $\alpha$, which we call Deep OSQP — Improved.
@@ 4.1 Deep FlexQP Architecture
Our proposed Deep FlexQP learns feedback policies for the algorithm parameters as a function of the current state of the optimizer as well as the QP and ADMM residuals, see Figure 2. Based on the successes of the Deep OSQP methods discussed above, we learn separate policies $\pi_{I}$, $\pi_{E}$, and $\pi_{\alpha}$ for the parameters related to the inequality constraints, equality constraints, and the relaxation parameter, respectively. Furthermore, the $\pi_{I}$ and $\pi_{E}$ policies are designed so that the resultant architecture is independent of problem size and permutation by applying the policies in a batched fashion per constraint coefficient. Note that the variables $s$, $z_{I}$, $w_{s}$ and $y_{I}$ are one-to-one with $\mu_{I}$, $\sigma_{s}$, and $\rho_{I}$, and that $z_{E}$ and $y_{E}$ are one-to-one with $\mu_{E}$ and $\rho_{E}$. We therefore use $s$, $z_{I}$, $w_{s}$, and $y_{I}$ as inputs to the inequality policy $\pi_{I}$, along with their associated ADMM residuals and the relaxed QP residual $\zeta_{I}=Gx+s-h-z_{I}$. We also include the infinity norm of the QP dual residual $\zeta_{\text{dual}}=Px+q+G^{\top}y_{I}+A^{\top}y_{E}$ as a scale-invariant measure of optimality. This leads to a total of ten inputs and three outputs corresponding to the coefficients of $\mu_{I}$, $\sigma_{s}$, and $\rho_{I}$. Likewise, the equality constraint policy $\pi_{E}$ predicts the coefficients of $\mu_{E}$ and $\rho_{E}$ as a function of the variables $z_{E}$ and $y_{E}$, along with their associated ADMM residuals, the relaxed QP residual $\zeta_{E}=Ax-b-z_{E}$, and the infinity norm of the dual residual $\zeta_{\text{dual}}$, leading to six total inputs. Finally, the policy $\pi_{\alpha}$ learns the relaxation parameter $\alpha$ as a function of the infinity norms of each of the QP and ADMM residuals, which provide a scale-invariant measure of how well and fast the optimizer is converging. Full expressions for the residuals and policies are given in Appendix F.
All policies are parameterized by long short-term memory (LSTM) networks (hochreiter1997long), with the hypothesis that learning long-term dependencies can aid the selection of the optimal parameters. This furthers the idea from saravanos2025deep that time-varying feedback on the current (nominal) parameters can provide a large improvement. In our case, we are applying feedback based on a latent state capturing the optimization history. Our results show that LSTMs provide the most benefit for problems where the active constraints might change many times over the course of the optimization. An ablation analysis and further discussion is provided in Appendix M.

### v2

@@ 3.3 Operator Splitting & ADMM
The optimization in Eq. 6 can be simplified further to make the terms appearing in the $\ell_{1}$ penalty functions easier to handle. Introducing decision variables $z_{I}\in{\mathbb{R}}^{m}$ and $z_{E}\in{\mathbb{R}}^{p}$, we have
$\operatorname*{minimize}_{x,s,z_{I},z_{E}}\quad$ $\frac{1}{2}x^{\top}Px+q^{\top}x+\mu_{I}\left\lVert z_{I}\right\rVert_{1}+\mu_{E}\left\lVert z_{E}\right\rVert_{1},$ (7a)
$\operatorname*{subject\ to}\quad$ $z_{I}=Gx+s-h,\hskip 5.0ptz_{E}=Ax-b,\hskip 5.0pts\geq 0.$ (7b)
The variables $z_{I}$ and $z_{E}$ are equal to the constraint violation, so their optimal values can be viewed as a certificate of feasibility for Eq. 3 if $z_{I}^{*}=z_{E}^{*}=0$ and infeasibility if $z_{I}^{*}\neq 0$ or $z_{E}^{*}\neq 0$. While it may seem tempting to apply ADMM to this formulation, the resultant updates will not have a closed-form solution no matter how the variable splitting is performed. Inspired by OSQP (Stellato et al., 2020), we perform a final transformation by introducing variables $\tilde{x}$, $\tilde{s}$, $\tilde{z}_{I}$, and $\tilde{z}_{E}$, yielding
$\operatorname*{minimize}_{\tilde{\bm{x}},\bm{x}}\quad$ $\underbrace{\frac{1}{2}\tilde{x}^{\top}P\tilde{x}+q^{\top}\tilde{x}+{\mathcal{I}}_{I}(\tilde{\bm{x}})+{\mathcal{I}}_{E}(\tilde{\bm{x}})}_{=\vcentcolon\,f(\tilde{\bm{x}})}+\underbrace{{\mathcal{I}}_{s}(\bm{x})+\mu_{I}\left\lVert z_{I}\right\rVert_{1}+\mu_{E}\left\lVert z_{E}\right\rVert_{1}}_{=\vcentcolon\,g(\bm{x})},$ (8a)
$\operatorname*{subject\ to}\quad$ $\tilde{\bm{x}}=\bm{x},$ (8b)
@@
${\mathcal{I}}_{I}(\bm{x})=\begin{cases}0&z_{I}=Gx+s-h,\\ +\infty&\text{otherwise},\end{cases}\hskip 5.0pt{\mathcal{I}}_{E}(\bm{x})=\begin{cases}0&z_{E}=Ax-b,\\ +\infty&\text{otherwise},\end{cases}\hskip 5.0pt{\mathcal{I}}_{s}(\bm{x})=\begin{cases}0&s\geq 0,\\ +\infty&\text{otherwise}.\end{cases}$
Let the dual variables for the constraint $\tilde{\bm{x}}=\bm{x}$ be $\bm{y}=(w_{x},w_{s},y_{I},y_{E})$. The ADMM updates for solving Eq. 8 are given by
$\tilde{\bm{x}}^{k+1}=\operatorname*{arg\,min}_{\tilde{\bm{x}}}\hskip 5.0ptf(\tilde{\bm{x}})+(\sigma_{x}/2)\lVert\tilde{x}-x^{k}+\sigma_{x}^{-1}w_{x}^{k}\rVert_{2}^{2}+(\sigma_{s}/2)\lVert\tilde{s}-s^{k}+\sigma_{s}^{-1}w_{s}^{k}\rVert_{2}^{2}$ (9) $+(\rho_{I}/2)\lVert\tilde{z}_{I}-z_{I}^{k}+\rho_{I}^{-1}y_{I}^{k}\rVert_{2}^{2}+(\rho_{E}/2)\lVert\tilde{z}_{E}-z_{E}^{k}+\rho_{E}^{-1}y_{E}^{k}\rVert_{2}^{2},$ $\hskip-7.5ptx^{k+1}=\alpha\tilde{x}^{k+1}+(1-\alpha)x^{k}+\sigma_{x}^{-1}w_{x}^{k},$ $s^{k+1}=\left(\alpha\tilde{s}^{k+1}+(1-\alpha)s^{k}+\sigma_{s}^{-1}w_{s}^{k}\right)_{+},$ $\hskip-7.5ptz_{I}^{k+1}=S_{\mu_{I}/\rho_{I}}(\alpha\tilde{z}_{I}^{k+1}+(1-\alpha)z_{I}^{k}+\rho_{I}^{-1}y_{I}^{k}),$ $z_{E}^{k+1}=S_{\mu_{E}/\rho_{E}}(\alpha\tilde{z}_{E}^{k+1}+(1-\alpha)z_{E}^{k}+\rho_{E}^{-1}y_{E}^{k}),$ $\hskip-7.5ptw_{x}^{k+1}=w_{x}^{k}+\sigma_{x}(\alpha\tilde{x}^{k+1}+(1-\alpha)x^{k}-x^{k+1}),$ $w_{s}^{k+1}=w_{s}^{k}+\sigma_{s}(\alpha\tilde{s}^{k+1}+(1-\alpha)s^{k}-s^{k+1}),$ $\hskip-7.5pty_{I}^{k+1}=y_{I}^{k}+\rho_{I}(\alpha\tilde{z}_{I}^{k+1}+(1-\alpha)z_{I}^{k}-z_{I}^{k+1}),$ $y_{E}^{k+1}=y_{E}^{k}+\rho_{E}(\alpha\tilde{z}_{E}^{k+1}+(1-\alpha)z_{E}^{k}-z_{E}^{k+1}),$
where $\sigma_{x},\sigma_{s},\rho_{I},\rho_{E}>0$ are the augmented Lagrangian penalty parameters, $\alpha\in(0,2)$ is the ADMM relaxation parameter, $(s)_{+}=\max(s,0)$ is the rectified linear unit (ReLU) activation function, and $S_{\kappa}(z)=(z-\kappa)_{+}-(-z-\kappa)_{+}$ is the soft thresholding operator, which is the proximal operator of the $\ell_{1}$ norm (Boyd et al., 2011). Since $x$ is unconstrained in the second block, we have that $w_{x}^{k+1}=0$ for all $k\geq 0$, so the $w_{x}$ variable and update can be disregarded. The first block update is the most computationally-demanding step of the algorithm and requires the solution of an equality-constrained QP. We show how to solve this QP using either a direct or indirect method in Appendix C; the indirect method becomes the only suitable choice for large-scale problems where the dimension can be very large. The final algorithm is summarized in Algorithm 1 of Appendix D.
@@ 3.4 Convergence & Theoretical Analysis
We establish the convergence of Algorithm 1 by showing that there always exists a saddle point of the Lagrangian for Eq. 8 for any $\mu_{I},\mu_{E}>0$, as long as the relaxed QP objective in Eq. 6 is not unbounded below. Then, since the optimization is a composite minimization of two closed, proper, convex functions, the algorithm converges by the general convergence of two-block ADMM (Boyd et al., 2011). The proof is provided in Appendix E.
@@
$\mathcal{L}(\tilde{\bm{x}},\bm{x},\bm{y})=f(\tilde{\bm{x}})+g(\bm{x})+\bm{y}^{\top}(\tilde{\bm{x}}-\bm{x}).$ (10)
The assumption that the objective is not unbounded below is called coercivity and is relatively weak (Bauschke and Combettes, 2017). Coercivity is only broken in the rare case when there exists an $x\neq 0$ such that $Px=0$, $Gx=0$, $Ax=0$, and $q^{\top}x<0$, which causes the optimal solution to diverge. In our cases of interest, we consider over-constrained problems with many more constraints than optimization variables, thus it is extremely unlikely that this assumption does not hold. Moreover, a bounded objective can be guaranteed if $P\succ 0$ or if $G$ or $A$ are full column rank.
The following theorem establishes the relationship between the FlexQP solution and the solution to the original QP and can be proven using the definition of soft thresholding. The proof is given in Appendix F.
@@
This shows that FlexQP solves Eq. 3 if the original QP is feasible, and otherwise identifies a stationary point of the infeasibilities, similar to Nocedal and Wright (2006, Theorem 17.4).
@@
Role of Elastic Penalty Parameters: The parameters $\mu_{I}$ and $\mu_{E}$ only appear in a single step of the algorithm as part of the soft thresholding in the second block ADMM updates. Larger elastic penalties $\mu$ result in a larger threshold, meaning that a larger amount of constraint violation will be zeroed out. The choice of $\mu$ is key for satisfying the conditions of Theorem 3.1.
@@
The parameter $\sigma_{s}$ plays the role of a quadratic cost on the slack variable $s$ when solving  Eq. 9. It also plays a small role in regularizing the constraint matrix $G$. In practice, tuning this parameter is the most difficult as the optimal value appears to depend strongly on the scaling of the objective and the constraints. This motivates our adaptive data-driven approach described in Section 4.
The penalty parameters $\rho_{I}$ and $\rho_{E}$ play two key roles in the algorithm. First, they regularize the constraint matrices $G$ and $A$ so that Eq. 9 is solvable regardless of the rank of $G$ or $A$. Second, they weight the noise level in the soft thresholding operations, playing an inverse role to $\mu_{I}$ and $\mu_{E}$, where a larger $\rho$ results in a smaller threshold. Determining the optimal values of these parameters by hand is unintuitive as they can have varying effects on the optimization, further motivating the deep unfolding approach presented in the next section.
@@ 4 Accelerating Quadratic Programming via Deep Unfolding
Figure 2: One layer of our proposed Deep FlexQP architecture. We learn dimension-agnostic feedback policies for the parameters while the propagation from one layer to the next is defined by the ADMM updates Eq. 9.
We focus our study on two recently proposed data-accelerated QP optimizers. The deep centralized QP optimizer from Saravanos et al. (2025) is a version of deep-unfolded OSQP where the penalty parameters $\rho$ and relaxation parameter $\alpha$ are learned as feedback policies on the problem residuals using an analogy to feedback control. In our comparisons, we refer to their method as Deep OSQP. The main limitation of their approach is that only scalar penalty parameters are learned, but it could be the case that different penalty parameters should be applied to different constraints to more effectively accelerate the optimizer. This was the main motivation for the deep QP method proposed by Ichnowski et al. (2021), where a policy that outputs a vector of penalty parameters is learned using reinforcement learning. The vector policy is applied across the constraint dimensions so it is dimension-agnostic and generalizes across different problem classes. The authors show that the vector policy outperforms the scalar one across a suite of QP benchmarks. However, unlike Saravanos et al. (2025), the authors do not learn the relaxation parameter $\alpha$, which can greatly improve the convergence of ADMM (Boyd et al., 2011). We implement the approach from Ichnowski et al. (2021) and train it using the supervised learning scheme from Saravanos et al. (2025), leading to the baseline Deep OSQP — RLQP Parameterization. Finally, we implement a best-of-both-worlds approach that learns a vector feedback policy for the penalty parameters $\rho$ while also learning a policy for the ADMM relaxation parameter $\alpha$, which we call Deep OSQP — Improved.
@@ 4.1 Deep FlexQP Architecture
Our proposed Deep FlexQP learns feedback policies for the algorithm parameters as a function of the current optimizer state as well as the QP and ADMM residuals (Fig. 2). We learn separate policies $\pi_{I}$, $\pi_{E}$, and $\pi_{\alpha}$ for the inequality constraints, equality constraints, and relaxation parameter, respectively. The constraint policies $\pi_{I}$ and $\pi_{E}$ take as input the ADMM variables along with the primal and dual QP residuals, and are applied in a batched fashion per constraint coefficient, making the architecture independent of problem size and permutation. The relaxation policy $\pi_{\alpha}$ takes as input the infinity norms of each residual, providing a scale-invariant representation of the convergence progress. Full expressions for the policy inputs and outputs are given in Appendix G.
All policies are parameterized by long short-term memory (LSTM) networks (Hochreiter and Schmidhuber, 1997), with the hypothesis that learning long-term dependencies can aid the selection of the optimal parameters. This furthers the idea from Saravanos et al. (2025) that time-varying feedback on the current (nominal) parameters can provide a large improvement. In our case, we are applying feedback based on a latent state capturing the optimization history. Our results show that LSTMs provide the most benefit for problems where the active constraints might change many times over the course of the optimization. An ablation analysis and further discussion is provided in Appendix N.